# Explainable post-training bias mitigation with distribution-based fairness metrics

### v2

@@ 14         Retrieve $f_{*}(X_{bias}),w(X_{bias})$
Constructing the efficient frontier of the family (4.1) amounts to solving the minimization problem (BM). Since any model from ${\cal F}(f_{*};w)$ is an adjustment of $f_{*}$ by a linear combination of encoders $\{w_{j}(\cdot;f_{*})\}_{j=0}^{m}$, we propose employing stochastic gradient descent, where the map $\theta\mapsto{\cal L}(\theta)+\omega\mathcal{B}(\theta)$ is approximated by appropriately designed differentiable estimators of bias metrics such as those in Section 3.
@@ 4.2.3 Explanation rebalancing
Note that the approaches we presented above are not mutually exclusive and may be combined without loss of explainability.
@@ 5 Experiments
We test the gradient-based methods we introduced in Section 4: additive model correction, tree rebalancing, and explanation rebalancing. For additive model correction, we treat all features as numerical and remove missing values by imputing the mean. For tree rebalancing, we employ forty weight values based on equation (4.5) with fixed coefficients $a_{kj}$ determined by PCA to avoid over-parametrization. For explanation rebalancing, we use CatBoost’s built-in marginal Shapley value implementation and thus refer to this particular methodology as Shapley rebalancing. Using the stochastic gradient descent strategy of algorithm 1, we minimize a cross-entropy loss subject to a fairness penalization based on (3.1). Specifically, the penalization uses the cost function $c(a,b)=h(a-b)=(a-b)^{2}$ that corresponds to the energy distance bias (3.10) and ensures that both the estimators (3.7) and (3.11) are unbiased. However, our results are still presented in the context of the Wasserstein metric (i.e. $h(z)=|z|$) due to its practical interpretation and relationship to prior literature (e.g. [44]). For further details about our gradient descent approach and custom loss function, see appendix A.4.
@@ 5.1 Synthetic datasets
The relative performances of methods on these synthetic datasets can be understood in the context of free parameters. Tree rebalancing has a free parameter for each of the forty principal components that rebalancing is applied to and optimal transport projection learns a new CatBoost model with a flexible number of parameters. In contrast, additive model correction, Shapley rebalancing, and predictor rescaling only have five parameters, one for each predictor. The similarities in the frontiers for additive model correction, Shapley rebalancing, and predictor rescaling are also no coincidence. When one is mitigating models linear in predictors, the model family explored by Shapley rebalancing is equivalent to the model family produced by predictor rescaling, and also equivalent to the model family produced by additive terms linear in the raw attributes. In our synthetic examples, the true score $f(x)$ is linear in predictors, so this correspondence is approximately true when mitigating trained models.
@@ 5.2 Real world datasets
Figure 2: Efficient frontiers for UCI Adult, UCI Bank Marketing, and COMPAS datasets. All results are presented on their respective test datasets.
@@
In contrast, UCI Bank Marketing has a limited number of observations from the protected class yet more features than UCI Adult. As a result, strategies susceptible to overfitting on $Y$ fare worse. Here, optimal transport projection performs the best as it utilizes a large number of free parameters without depending on $Y$, which it may otherwise overfit to. This method is followed by additive model correction, predictor rescaling, and Shapley rebalancing. These methods perform similarly and directly use $Y$ in the optimization procedure. However, they are still less susceptible to overfitting than the least optimal strategy on this dataset: tree rebalancing, which has the greatest number of free parameters and directly optimizes using $Y$.
@@
5.3 Addressing overfitting in SGD methods
In section 5.2, we saw some evidence that mitigation methods employing high dimensional optimization may fail when they can overfit to the response $Y$. For example, SGD based methods like tree rebalancing and Shapley rebalancing with access to $Y$ struggle on UCI Bank Marketing while other high-dimensional methods, such as optimal transport projection without access to $Y$, thrive. In such cases, removing the explicit use of $Y$ in BM may improve test dataset performance. In the classification setting, we propose substituting ${\cal L}(\theta)=\mathbb{E}[L(\sigma\circ f(X;\theta),Y)]$ with an alternative loss term $\mathbb{E}[\tilde{L}(\sigma\circ f(X;\theta),\sigma\circ f_{*}(X))]$ with $\tilde{L}$ given as follows:
$\tilde{L}(p,q):$ $=p\log\left(\frac{p}{q}\right)+(1-p)\log\left(\frac{1-p}{1-q}\right)=D_{KL}(p||q)+D_{KL}(1-p||1-q)\,$,
where $D_{KL}$ is the Kullback–Leibler divergence and $\tilde{L}$ may be interpreted as a cross-entropy.
Figure 3: Efficient frontiers for the UCI Bank Marketing dataset using log loss vs Wasserstein-1 bias and AUC vs KS bias with additive model correction, Shapley rebalancing, and tree rebalancing methods to training without direct use of the response variable.
Applying additive model correction, tree rebalancing, and Shapley rebalancing mitigation methods to the UCI Bank Marketing dataset with this new modified loss yields Figure 3, which compares these updated methods with previously shown frontiers. With the $Y$-unaware loss function, tree rebalancing and Shapley rebalancing beat predictor rescaling across multiple bias-performance metric pairs and are nearly at the level of optimal transport projection on UCI Bank Marketing. The performance of the additive model correction approach is mostly unchanged and is similar to the new tree rebalancing and Shapley rebalancing frontiers. Perhaps due to the simplicity of using raw features, this method had less of an issue with fitting to noise.
Furthermore, compared to Figure 2, the frontiers for tree rebalancing and Shapley rebalancing are considerably improved in absolute terms. This exercise demonstrates that, while SGD mitigation methods can overfit due to their high-dimensionality, the framework is sufficiently flexible to allow for solutions.
Acknowledgements
The authors thank Alex Lin (Lead Research Scientist, DFS) for his valuable comments and editorial suggestions that aided us in writing this article. We also thank Arjun Ravi Kannan (Director, Modeling, DFS) and Stoyan Vlaikov (VP, Data Science, DFS) for their helpful business and compliance insights.
@@ A.1 Datasets
UCI Adult: The UCI Adult dataset includes five numerical variables (Age, education-num, capital gain, capital loss, hours per week) and seven categorical variables (workclass, education, marital-status, occupation, relationship, race, native-country) for a total of twelve independent variables. In addition, UCI Adult includes a numerical dependent variable (income) and gender information (male and female) for each record. We encode the categorical variables with ordinal encoding and we binarize income based on whether it exceeds \$50,000.
@@
UCI Bank Marketing: The UCI Bank Marketing dataset includes thirteen numerical variables (default, housing, loan, duration, campaign, pdays, previous, poutcome, emp.var.rate, cons.price.idx, cons.conf.idx, euribor3m, nr.employed) and six categorical variables (education, month, day_of_week, job, marital, contact). We convert education to a numerical variable based on the length of schooling suggested by its categories. Similarly, we convert month and day_of_week to numerical variables with month going from zero to eleven (January through December) and day_of_week going from zero to four (Monday through Friday). We encode the rest of the categorical variables using ordinal encoding. Furthermore, we represent categories like unknown or non-existent as missing to leverage CatBoost’s internal handling of missing values. The dependent variable is a yes/no classification reflecting whether marketing calls to a client yielded a subscription. UCI Adult also includes age information which is treated as sensitive demographic information.
@@
COMPAS: The COMPAS dataset includes three numerical variables (priors_count, two_year_recid) and three categorical variables (c_charge_degree, sex, age). Age is encoded from zero to two in order of youngest to oldest age group. Sex and c_charge_degree are ordinal encoded. The dependent variable describes the risk classification of the COMPAS algorithm, which we binarize as zero if low and one otherwise. COMPAS also includes race information which is treated as sensitive demographic information. Following Zafar et al. [70], we only include records when days_b_screening_arrest $\in[-30,30]$, is_recid $\neq-1$, c_charge_degree $\neq 0$ and the risk score is available.
@@ A.3 Predictor rescaling methodology
Predictor rescaling as described in Miroshnikov et al. [43] was implemented with $n_{prior}=100$, $n_{bo}=50$, and $((\omega_{j}))_{j=1}^{21}=((0.5\cdot j-0.5)_{j=1}^{21}$; see Algorithm 2 (which originates from that paper). Bayesian optimization parameters were given by $\kappa=1.5$ and $\xi=0.0$. We also implement predictor rescaling using $1150$ iterations of random search. The results of predictor rescaling we present combine the best of both these methods. In this work, we make use of a linear compressive family with transformations
@@
  • (UCI Adult) The top five most biased predictors: relationship (1st), marital-status (2nd), hours per week (3rd), Age (4th), capital gain (5th)
  • (UCI Bank Marketing) The top eight most biased predictors —nr.employed (1st), euribor3m (2nd), month (3rd), cons.price.idx (4th), duration (5th), emp.var.rate (6th), pdays (7th) , cons.conf.idx (8th)
  • (COMPAS) All predictors: priors_count (1st), age (2nd), two_year_recid (3rd), c_charge_degree (4th), sex (5th)
@@ A.7 Bias mitigation using neural networks
Figure 5: Efficient frontiers for UCI Adult, UCI Bank Marketing, and COMPAS datasets. All results are presented on their respective test datasets.
@@
In contrast, neural networks are better positioned on the UCI Bank Marketing dataset. As noted in Section 5.3, this dataset is much more prone to overfitting than Adult due to fewer data points and a greater number of features. Consequently, the performance gap between more expressive models (i.e. CatBoost) and less expressive models (i.e., the neural network architectures under consideration) is smaller. As a result, the neural network frontier bests several frontiers yielded by the post-processing methods. Furthermore, perhaps because linear models are more prone to score compression than non-linear models, neural networks eclipse all gradient descent based post-processing methods on the distribution invariant metric pair (AUC vs. KS).
@@ Appendix C Model bias with multiple protected attributes
In this section, following the formulation in [44], we generalize the notion of model bias to the case of protected attributes with multiple classes.
To this end, let $f$ be a model, $X=(X_{1},\dots,X_{n})$ the predictors, and $G\in\{0,\dots,K-1\}$ a protected attribute, with $G=0$ denoting the non-protected (majority) class. Suppose that $\text{\rm Bias}(f|X,G\in\{k,m\})$ is a model bias metric that estimates the distance (or divergence) between any two subpopulation distributions $P_{f(X)|G=k}$ and $P_{f(X)|G=m}$. In that case, the total weighted model bias is defined to be
@@
It is worth noting that for $q>1$, there exists a unique minimizer of the right-hand side of (C.1), while for $q=1$, uniqueness fails. For this reason, $\mu_{B}^{(1)}$ can be obtained as the limit of $\mu_{B}^{(q)}$ as $q\to 1^{+}$; for details, see [7, Theorem 5.1].
@@ Appendix D Relaxation of distributions
Suppose $\lim_{s\to\infty}r_{s}(z)=H(z)$ for all $z\in\mathbb{R}$. Since $c$ is continuous on $0,1]^{2}$, by Lemma D.1(ii)(ii), we have $\lim_{s\to\infty}c(F^{(s)}_{Z_{0}}(t),F^{(s)}_{Z_{1}}(t))=c(F_{Z_{0}}(t),F_{Z_{1}}(t))$ for all $t\in\mathbb{R}$. Then, since $c$ is bounded on $0,1]^{2}$, by the dominated convergence theorem, we obtain (D.2). This gives $(i)$.
@@
Fix $s>0$. Clearly, to show that $\hat{I}_{m_{0},m_{1},T}^{(s)}:=\frac{1}{T}\sum_{j=1}^{T}h(\hat{\bf B}_{s,m_{0},m_{1}}({\bf t}^{(j)}))$ converges in the mean squared sense to $\int h(B_{s}(t))\,\mu(dt)=\mathbb{E}_{t\sim\mu}[h(B_{s}(t))]$, we require
@@ Appendix E Quantile transformed distributions with atoms
Thus, (E.7) allows for the generalization of Becker et al. [3, Theorem 3.4] for the case when the classification scores have atoms; see Proposition 2.1 in the main text. We note that a similar adjustment as in (2.5) is required for other types of bias such as Equal Opportunity ($\rm EO$}) and Predictive Equality ($\rm PE$}), as discussed in Theorem 3.4 of [3].
@@
The distribution-invariant model bias (2.4), assuming $P_{Z}$ has a density, can be expressed as follows [3, Theorem 3.4]:
@@ Appendix F Estimators for distribution-invariant bias metrics
Estimator 1: Monte Carlo. Here, we derive an analogue of the MC estimator (3.5) of the relaxed bias metric (3.4) in the case where $\mu=P_{f(X)}$.
@@
Estimator 2: Threshold-discrete. Next, we derive an analogue of the threshold-discrete estimator (3.7) of the relaxed bias metric (3.4) in the case where $\mu=P_{f(X)}$ has a density.
@@
$\int_{0}^{1}h\big{(}B_{s}(t;\theta)\big{)}\mu_{\theta}(dt)=\int_{0}^{1}h\big{(}B_{s}(t;\theta)\big{)}\rho_{\theta}(t)dt\approx\Big{(}\sum_{j=1}^{T}h(\hat{B}_{s}(t^{(j)};\theta))\rho(t_{j};\theta)\Delta t\Big{)}.$
@@
Estimator 3: Energy. Finally, consider the estimator (3.11) of the bias metric in (3.10), where $h(\cdot)=2|\cdot|^{2}$, $\mu$ is replaced with $\mu_{\theta}$, and $f(\cdot;\theta)$, $D_{k}$, and $D_{f(\cdot;\theta)}$ are as defined above.

### v3

@@ 14         Retrieve $f_{*}(X_{bias}),w(X_{bias})$
Constructing the efficient frontier of the family (4.1) amounts to solving the minimization problem (BM). Since any model from ${\cal F}(f_{*};w)$ is an adjustment of $f_{*}$ by a linear combination of encoders $\{w_{j}(\cdot;f_{*})\}_{j=0}^{m}$, we propose employing stochastic gradient descent, as outlined in Algorithm 1, where the map $\theta\mapsto{\cal L}(\theta)+\omega\mathcal{B}(\theta)$ is approximated by appropriately designed differentiable estimators of bias metrics such as those in Section 3.
@@ 4.2.3 Explanation rebalancing
The approaches we presented above are not mutually exclusive and may be combined without affecting explainability.
@@ 5 Experiments
We test the gradient-based methods we introduced in Section 4: additive model correction, tree rebalancing, and explanation rebalancing. For additive model correction, we treat all features as numerical and remove missing values by imputing the mean. For tree rebalancing, we employ forty weight values based on equation (4.5) with fixed coefficients $a_{kj}$ determined by PCA to avoid over-parametrization. For explanation rebalancing, we use CatBoost’s built-in marginal Shapley value implementation and thus refer to this particular methodology as Shapley rebalancing. Using the stochastic gradient descent strategy of algorithm 1, we minimize a cross-entropy loss subject to a fairness penalization based on (3.1). Specifically, the penalization uses the cost function $c(a,b)=h(a-b)=(a-b)^{2}$ that corresponds to the energy distance bias (3.10) and ensures that both the estimators (3.7) and (3.11) are unbiased. However, our results are still presented in the context of the Wasserstein metric (i.e. $h(\cdot)=|\cdot|$) due to its practical interpretation and relationship to prior literature (e.g. [20]). For further details on our gradient descent approach and custom loss function, see Appendix A.4.
@@ 5.1 Synthetic datasets
The relative performances of methods on these synthetic datasets can be understood in the context of free parameters. Tree rebalancing has a free parameter for each of the forty principal components that rebalancing is applied to and optimal transport projection learns a new CatBoost model with a flexible number of parameters. In contrast, additive model correction, Shapley rebalancing, and predictor rescaling only have five parameters, one for each predictor. The similarities in the frontiers for additive model correction, Shapley rebalancing, and predictor rescaling are also no coincidence. When one is mitigating models linear in predictors, the model family explored by Shapley rebalancing is equivalent to the model family produced by predictor rescaling, and also equivalent to the model family produced by additive terms linear in the raw attributes assuming no limitations on mitigation parameters. In our synthetic examples, the true score $f(x)$ is linear in predictors, so this correspondence is approximately true when mitigating trained models.
@@ 5.2 Real world datasets
Figure 2: Efficient frontiers for UCI Adult, UCI Bank Marketing, and COMPAS datasets, evaluated on their test datasets.
@@
In contrast, UCI Bank Marketing has a limited number of observations from the protected class yet more features than UCI Adult. Here, optimal transport projection outperforms the rest of the post-processing methods considerably and remaining differences in frontiers are fairly minor. Since optimal transport projection is the only method that does not explicitly depend on $Y$, its performance may be due to a lack of overfitting relative to the other methods.
@@
Remark 5.1 (Addressing overfitting in SGD methods).
Mitigation methods employing high-dimensional optimization may overfit to the response variable $Y$. In such cases, removing the explicit use of $Y$ in (BM) may improve test dataset performance. Following Hinton’s work on knowledge distillation in the classification setting [66], we propose substituting ${\cal L}(\theta)=\mathbb{E}[L(\sigma\circ f(X;\theta),Y)]$ with an alternative loss term $\mathbb{E}[\tilde{L}(\sigma\circ f(X;\theta),\sigma\circ f_{*}(X))]$, where $\tilde{L}$ is defined as:
$\tilde{L}(p,q):$ $=D_{KL}(\text{Bern}(p)||\text{Bern}(q))=p\log\left(\frac{p}{q}\right)+(1-p)\log\left(\frac{1-p}{1-q}\right)\,$,
where $D_{KL}$ denotes the Kullback–Leibler divergence and $\text{Bern}(p)$ the Bernoulli distribution.
@@ A.1 Datasets
UCI Adult.
The UCI Adult dataset includes five numerical variables (Age, education-num, capital gain, capital loss, hours per week) and seven categorical variables (workclass, education, marital-status, occupation, relationship, race, native-country) for a total of twelve independent variables. In addition, UCI Adult includes a numerical dependent variable (income) and gender information (male and female) for each record. We encode the categorical variables with ordinal encoding and we binarize income based on whether it exceeds \$50,000.
@@
UCI Bank Marketing.
The UCI Bank Marketing dataset includes thirteen numerical variables (default, housing, loan, duration, campaign, pdays, previous, poutcome, emp.var.rate, cons.price.idx, cons.conf.idx, euribor3m, nr.employed) and six categorical variables (education, month, day_of_week, job, marital, contact). We convert education to a numerical variable based on the length of schooling suggested by its categories. Similarly, we convert month and day_of_week to numerical variables with month going from zero to eleven (January through December) and day_of_week going from zero to four (Monday through Friday). We encode the rest of the categorical variables using ordinal encoding. Furthermore, we represent categories like unknown or non-existent as missing to leverage CatBoost’s internal handling of missing values. The dependent variable is a yes/no classification reflecting whether marketing calls to a client yielded a subscription. UCI Adult also includes age information which is treated as sensitive demographic information.
@@
COMPAS.
The COMPAS dataset includes three numerical variables (priors_count, two_year_recid) and three categorical variables (c_charge_degree, sex, age). Age is encoded from zero to two in order of youngest to oldest age group. Sex and c_charge_degree are ordinal encoded. The dependent variable describes the risk classification of the COMPAS algorithm, which we binarize as zero if low and one otherwise. COMPAS also includes race information which is treated as sensitive demographic information. Following Zafar et al. [65], we only include records when days_b_screening_arrest $\in[-30,30]$, is_recid $\neq-1$, c_charge_degree $\neq 0$ and the risk score is available.
@@ A.3 Predictor rescaling methodology
Predictor rescaling as described in Miroshnikov et al. [21] was implemented with $n_{prior}=100$, $n_{bo}=50$, and $\{\omega_{j}\}_{j=1}^{21}=\{\tfrac{1}{2}\cdot(j-1)\}_{j=1}^{21}$; see Algorithm 2 below, which originates from that paper. Bayesian optimization parameters were given by $\kappa=1.5$ and $\xi=0.0$. We also implement predictor rescaling using $1150$ iterations of random search. The results of predictor rescaling we present combine the best of both these methods. In this work, we make use of a linear compressive family with transformations
@@
  • $\bullet$
    (UCI Adult) The top five most biased predictors: relationship (1st), marital-status (2nd), hours per week (3rd), Age (4th), capital gain (5th).
  • $\bullet$
    (UCI Bank Marketing) The top eight most biased predictors —nr.employed (1st), euribor3m (2nd), month (3rd), cons.price.idx (4th), duration (5th), emp.var.rate (6th), pdays (7th) , cons.conf.idx (8th).
  • $\bullet$
    (COMPAS) All predictors: priors_count (1st), age (2nd), two_year_recid (3rd), c_charge_degree (4th), sex (5th).
@@ A.6 Impact of dimensionality on predictor rescaling methods
Figure 4: Efficient frontiers for UCI Adult, Bank Marketing, and COMPAS datasets, evaluated on their test datasets.
@@ A.7 Bias mitigation using neural networks
In contrast, neural networks are better positioned on the UCI Bank Marketing dataset, achieving a best performance value lower than, but much closer to, that of the trained (CatBoost) base model. As a result, the neural network frontier bests several others (tree and Shapley rebalancing) outside the high performance regime. Furthermore, perhaps because linear models are more prone to score compression than non-linear models, neural networks beat all gradient descent based post-processing methods on the distribution invariant metric pair (AUC vs. KS) in the low bias regime.
@@
For more context on methodology, bias mitigation for neural networks was conducted using the Karush-Kuhn-Tucker approach [39, 40] in a manner analogous to that used by our post-processing methodologies (using an adaptation of Algorithm 1 for parameters of non-linear models). As in A.4, we use minimization objectives of the form
$(1-\omega_{j}){\cal L}+\omega_{j}\mathcal{B},$
@@ Appendix C Model bias with multiple protected attributes
In this section, following the formulation in [20], we generalize the notion of model bias to the case of protected attributes with multiple classes. To this end, let $f$ be a model, $X=(X_{1},\dots,X_{n})$ the predictors, and $G\in\{0,\dots,K-1\}$ a protected attribute, with $G=0$ denoting the non-protected (majority) class. Suppose that $\text{\rm Bias}(f|X,G\in\{k,m\})$ is a model bias metric that estimates the distance (or divergence) between any two subpopulation distributions $P_{f(X)|G=k}$ and $P_{f(X)|G=m}$. In that case, the total weighted model bias is defined to be
@@
It is worth noting that for $q>1$, there exists a unique minimizer of the right-hand side of (C.1), while for $q=1$, uniqueness fails. For this reason, $\mu_{B}^{(1)}$ can be obtained as the limit of $\mu_{B}^{(q)}$ as $q\to 1^{+}$; for details, see [17, Theorem 5.1].
@@ Appendix D Relaxation of distributions
Suppose $\lim_{s\to\infty}r_{s}(z)=H(z)$ for all $z\in\mathbb{R}$. Since $c$ is continuous on $0,1]^{2}$, by Lemma D.1(ii)(ii), we have
$\lim_{s\to\infty}c(F^{(s)}_{Z_{0}}(t),F^{(s)}_{Z_{1}}(t))=c(F_{Z_{0}}(t),F_{Z_{1}}(t))$
for all $t\in\mathbb{R}$. Then, since $c$ is bounded on $0,1]^{2}$, by the dominated convergence theorem, we obtain (D.2). This gives $(i)$.
@@
Fix $s>0$. To show that
$\hat{I}_{m_{0},m_{1},T}^{(s)}:=\frac{1}{T}\sum_{j=1}^{T}h(\hat{\bf B}_{s,m_{0},m_{1}}({\bf t}^{(j)}))$
converges in the mean squared sense to
$\int h(B_{s}(t))\,\mu(dt)=\mathbb{E}_{t\sim\mu}[h(B_{s}(t))],$
we require
@@ Appendix E Quantile transformed distributions with atoms
Thus, (E.7) allows for the generalization of Becker et al. [24, Theorem 3.4] for the case when the classification scores have atoms; see Proposition 2.1 in the main text. We note that a similar adjustment as in (2.5) is required for other types of bias such as Equal Opportunity ($\rm EO$}) and Predictive Equality ($\rm PE$}), as discussed in Theorem 3.4 of [24].
@@
The distribution-invariant model bias (2.4), assuming $P_{Z}$ has a density, can be expressed as follows [24, Theorem 3.4]:
@@ Appendix F Estimators for distribution-invariant bias metrics
Estimator 1: Monte Carlo.
Here, we derive an analogue of the MC estimator (3.5) of the relaxed bias metric (3.4) in the case where $\mu=P_{f(X)}$.
@@
Estimator 2: Threshold-discrete.
Next, we derive an analogue of the threshold-discrete estimator (3.7) of the relaxed bias metric (3.4) in the case where $\mu=P_{f(X)}$ has a density.
@@
$\int_{0}^{1}h\big(B_{s}(t;\theta)\big)\mu_{\theta}(dt)=\int_{0}^{1}h\big(B_{s}(t;\theta)\big)\rho_{\theta}(t)dt\approx\Big(\frac{1}{T}\sum_{j=1}^{T}h(\hat{B}_{s}(t^{(j)};\theta))\rho(t_{j};\theta)\Big).$
@@
Estimator 3: Energy.
Finally, consider the estimator (3.11) of the bias metric in (3.10), where $h(\cdot)=2|\cdot|^{2}$, $\mu$ is replaced with $\mu_{\theta}$, and $f(\cdot;\theta)$, $D_{k}$, and $D_{f(\cdot;\theta)}$ are as defined above.
@@
Acknowledgements
The authors thank Alex Lin (Lead Research Scientist, Discover) for his valuable comments and editorial suggestions that aided us in writing this article. We also thank Arjun Ravi Kannan (Director, Modeling, Discover) and Stoyan Vlaikov (VP, Data Science, Discover) for their helpful business and compliance insights.
Authors’ contributions
Ryan Franks designed the high-level gradient-descent-based approach for optimizing encoder weights in post-processed models, developed the modified explainable variant of the optimal transport projection used for comparison, and proposed the use of energy distances and related unbiased estimators for bias mitigation. He also designed and implemented all experiments on public datasets and drafted the manuscript.
Alexey Miroshnikov proposed the conceptual framework for bias mitigation in ML models through post-processing methods optimized with a custom loss penalized by distribution-based fairness metrics. He introduced new fairness metrics, analyzed their theoretical properties, and developed differentiable estimators for their stochastic-gradient-descent optimization. He also proposed the use of weak learners as encoders in fairness optimization and contributed to revising the manuscript.
Kostas Kotsiopoulos performed the asymptotic analysis of differentiable estimators for non-invariant fairness metrics and provided technical proofs. He also proposed consistent estimators for a broad class of distribution-invariant fairness metrics and contributed to revising and editing the manuscript.